OpenStudy (anonymous):

The table shows the solution to the equation |2x - 4| - 3 = 3: Step 1 |2x - 4| = -3 + 3 Step 2 |2x - 4| = 0 Step 3 2x - 4 = 0 Step 4 2x = 4 Step 5 x = 2 Which is the first incorrect step? Step 1 Step 2 Step 3 Solution is correct

OpenStudy (anonymous):

|2x-4|-3=3 |2x-4|=3+3 (here -3 is incorrect it should not be negative |2x-4|=6

OpenStudy (anonymous):

on removing mod \[2x-4=\pm6\]

OpenStudy (anonymous):

2x-4=6 and 2x-4=-6 2x=4+6=10 2x=-6+4=-2 x=5 x=-1 x=5 or x=-1

OpenStudy (anonymous):

step 2 was wrong?

OpenStudy (anonymous):

we can see step 1 is incorrect

OpenStudy (anonymous):

loll step was ur Q in my ans

OpenStudy (anonymous):

can you help me with 2 more

OpenStudy (anonymous):

The table below shows two equations: Equation 1 |2x - 3| + 5 = 4 Equation 2 |5x + 3| - 10 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = 2, -3.2. The solutions to equation 1 are x = 1, 2 and equation 2 has no solution. The solutions to equation 1 are x = 1, 2 and equation 2 has solutions x = 2, -3.2.

OpenStudy (anonymous):

2is true